# Operating System And Artificial Intelligence: A Systematic Review

### v1

@@ 7.4 Findings
In essence, these advanced OS architectures elevate performance and security, fostering an innovative landscape for AI4OS. By catering to AI application requirements, they promise to rewrite the rules of computing, blending AI and OS synergies for smarter, more efficient, and secure computing futures.
8 RQ5: How LLMs create opportunities for OS?
The integration of LLMs into OSes presents a significant opportunity to enhance the user experience and overall system functionality[67][68]. LLMs, with their advanced natural language processing capabilities,[62][63][64] can transform the way users interact with and manage their computing environments.[76] We simply divide these works into two big categories: LLM AS OS and LLM4OS.
8.1 LLM AS OS
”LLM AS OS” refers to the integration of LLMs into the core of an OS, effectively serving as the ”brain” of the system,make OS capable of understanding and responding to natural language commands, thereby enabling more intuitive and flexible human-computer interaction[70][69].
Kamath et al.’s LLaMaS[54] addresses the complexities of diverse computing environments by leveraging LLMs to ease OS challenges in hardware adaptation and resource management. It adapts to new devices by interpreting plain text specs, recommending optimized OS tactics. The LLM frontend deciphers system characteristics, converting them into actionable embeddings. The backend uses these for on-the-fly OS decisions, like memory tier data movement. An experiment with ChatGPT demonstrated its skill in adjusting memory allocation for CPU and GPU tasks based on usage. Aimed at reducing administrative and research costs, LLaMaS autonomously aligns with hardware changes via LLMs’ innate zero-shot learning, eliminating the need for manual adjustments or detailed device-specific coding.
AIOS-Agent ecosystem[55][56], where LLMs act essentially as an ”operating system with soul”.LLMs are embedded in the OS kernel for intelligent decision-making and resource allocation. Its context window acts as memory, managing relevant data, while external storage serves as a file system with enhanced retrieval capabilities. Hardware and software are treated as peripherals and libraries, enabling agent-environment interaction. Natural language becomes the primary programming interface, democratizing software development. This ecosystem supports single and multi-agent applications for executing a broad range of tasks.
MemGPT[57] presents a solution to overcome the fixed-length context window limitation in LLMs, a hurdle for tasks demanding deep analysis of lengthy dialogues or extensive texts. Drawing parallels with traditional OS hierarchical memory systems, MemGPT introduces virtual context management, echoing virtual memory principles. It enables LLMs to access information beyond immediate capacity, mirroring OS memory management through strategic ’paging’. This system reacts to events like user messages, system alerts, and timers, appending pertinent data to the primary context buffer before processing. Additionally, it supports function chaining for uninterrupted, sequential task execution, accommodating intricate operations and long-term planning within the LLM’s context limitations.
8.2 LLM4OS
SecRepair[58], designed to address the challenge of identifying and repairing code vulnerabilities in software development. This system is powered by a LLM and incorporates reinforcement learning and semantic rewards to enhance its capabilities.To support the training process and prepare a robust model for vulnerability analysis, the authors have compiled and released a comprehensive instruction-based vulnerability dataset.They also propose a reinforcement learning technique with a semantic reward mechanism to generate concise and appropriate code commit comments. This technique is inspired by how humans fix code issues and provides developers with a clear understanding of the vulnerability and its resolution.
Rahman et al. [71] introduce ChronoCTI, an automated pipeline for mining temporal dynamics between attacker actions from text-based accounts of cyber incidents. The focus is on pinpointing repetitive action sequences, termed temporal attack patterns, crucial for preemptive defense strategies by security professionals against impending cyber threats. ChronoCTI is anchored by a curated dataset linking sentences to adversary maneuvers, temporal linkages across 94 attack narratives, and large language models refined on cybersecurity topics. The research delineates a structured approach, harnessing cutting-edge language models, NLP methods, and ML techniques to decipher the temporal structure within attack stories.
8.3 Findings
Addressing Research Question 5, our analysis highlights the incorporation of LLMs within operating systems heralds a transformative era, reshaping system dynamics and enhancing user engagement. They can enrich user interfaces by facilitating natural language processing for voice commands and text inputs, leading to more intuitive and personalized user experiences. In the realm of development, LLMs can automate code generation and optimization, streamlining coding processes and reducing errors. Additionally, they contribute to dynamic system management by analyzing usage patterns to optimize resource allocation, boosting system responsiveness. LLMs also play a pivotal role in intelligent troubleshooting, diagnosing system issues more accurately and recommending preventive measures to minimize downtime. Furthermore, they bolster OS security by detecting anomalies indicative of malicious activities and assisting in secure configuration management. Lastly, LLMs act as intelligent aids for developers, offering insights into function queries, suggesting best practices, and even generating documentation, thus accelerating the development process and enhancing code quality. Through continuous learning, LLMs ensure that OS performance and efficiency improve over time, adapting to user needs and operational contexts.
@@ 9 Opportunities for Future Work
9.1 Optimized OS for AI
The escalating sophistication of AI workloads, characterized by high-velocity I/O, minimal latency, and bandwidth-intensive requirements, is catalyzing an urgent need for AI-specialized OS optimization. As AI permeates diverse sectors, the imperative for a finely calibrated infrastructure escalates, poised to support the burgeoning demands of next-generation applications. The emergence of AI-dedicated hardware has dramatically reshaped AI computation landscapes. These technological marvels, while delivering unrivaled performance, necessitate software ecosystems meticulously attuned to their singular attributes.
In this multifaceted hardware ecosystem [20][23], the need for an OS that can seamlessly manage and orchestrate AI tasks across a spectrum of devices is more pressing than ever. AI has transcended beyond the realm of traditional neural networks, with innovative models leading the change in natural language processing and graph neural networks adeptly handling intricate relational data. These advanced applications are calling for specialized optimizations that a one-size-fits-all approach cannot provide. Juggling performance, energy conservation, and scalability constitutes a multifaceted challenge, demanding an OS equipped for dynamic adjustment to cater to the singular requisites of every workload.
Portatble OS for AI
In industrial settings, IoT devices generate massive amounts of data. Deploying deep learning models directly on edge devices (such as sensors, gateways, or edge servers) in equally important as relying solely on cloud-based processing. In a certain sense, deploying on edge devices can be a more challenging scenario because the computational power and bandwidth, of edge devices are more constrained[77][12][1]. Due to the diversity of edge devices,current AI platforms’ lack of portability across different edge platforms hinders widespread adoption ,and bypassing the kernel and directly accessing hardware while adapting to various hardware configurations can be extremely challenging. Therefore, an operating system that can accommodate multiple hardware types and is specifically tailored for AI applications is a potential solution[7][8].
Fast deployment and standard API interface
Simplified deployment means that even non-experts can utilize AI applications effectively. Developers, data scientists, and industrial engineers can quickly integrate the framework into their existing edge devices without extensive configuration or manual setup. And a standard API interface ensures that software components can communicate seamlessly with each other.
@@ 9.1 Optimized OS for AI
Resource-constrained environments, such as edge devices or embedded systems, often operate with tight constraints on memory, processing power, and energy while in scenarios like cloud computing or data centers, minimizing resource consumption directly impacts operational expenses. As OS and applications share the resources, the OS consumption determines the actual operational performance of the algorithmic model. So there is a need to design an efficient OS that can achieve small memory and fast training and inference. There are several promising approaches to achieve this. Apart from bypassing the kernel, unikernel OSes are designed to be fast and lightweight[11][32].
@@
The problem of privacy preservation in the context of machine learning is quite different from that in traditional data privacy protection, as machine learning can act as both friend and foe. Increasingly, edge devices are equipped with AI applications. This trend comes with privacy risks as models can leak information about their training data [24]. OS need to find ways to provide safeguards for ensuring the confidentiality and integrity of its data and programs.
9.2 Intelligent OS
9.2.1 Potential for Intelligent LibOS
@@
A recently emerging trend of Internet-based software systems is “resource adaptive,” i.e., software systems should be robust and intelligent enough to the changes of heterogeneous resources[9], both physical and logical, provided by their running environment[26]. The key principle of such an OS is based on resource disaggregation, resource provisioning as a service, and learning-based resource scheduling and allocation. Meanwhile, the OS should leverage advanced machine/deep learning techniques to derive configurations and policies and automatically learn to tune itself and schedule resources. One general and efficient Spark tuning framework is proposed and applied it to Tencent’s data platform[5].
@@
Microservices excels notably in granular security, isolated updates, and performance optimization through dynamic service management. AI technologies have proven adept at governing microservices, enabling intelligent service orchestration and customization, as highlighted in research [65] and [66], setting the stage for advanced system control and optimization[81].LibOS similarly exhibit gains in agility, composability, enhanced isolation, and facilitated deployment across heterogeneous hardware. LibOS presents a more auspicious ground for intelligent OS. There are at least 5 ways where LibOS will benefit to enable a Intelligent OS:
@@
FlexOS[82] presents a modular LibOS framework, comprised of discrete elements that can be segregated using a spectrum of hardware safeguards, complemented by adaptable data sharing protocols and software fortification methods. FlexOS adopts micro-libraries as fundamental units for segmentation. Integration or isolation of these components is dynamically managed, guided by safety and performance requirements.
Drawing from the microservices paradigm’s strengths in isolation and precise access governance, we posit that LibOS offers a strategic avenue to uphold data confidentiality during processing and safeguard overall system integrity. In the AI domain, where security and privacy are paramount[37]. LibOS’s robust isolation features make them ideally suited for secure operation in environments with concurrent tasks of varying privacy sensitivity. The innate simplicity and modularity of LibOS create an ecosystem that effectively quarantines critical workloads, bolstering confidence and aligning with rigorous privacy statutes.
Heterogeneous deployment
In the IoT epoch, a surge in varied endpoint devices underscores the imperative for AI integration. Yet, conventional monolithic OS kernels struggle, weighed down by excessive resource demands and challenges in hardware acclimatization, hindering their agility in today’s rapidly transforming milieu. In alignment, LibOS surface as a beacon of hope for versatile, cross-platform deployment amid the AI revolution. The Demikernel project[52] substantiates LibOS’s flexibility, enabling applications to operate fluidly across disparate devices sans customization. In LibOS’s schema, hardware interface components are architected as pluggable modules, amplifying interoperability and malleability.The intrinsic modularity and adaptability simplify AI assimilation into a spectrum of devices, nurturing a more dynamic and reactive IoT framework.
@@
A significant advancement brought about by the microservices paradigm is its capacity for orchestration for components,exemplified by tools like ISTIO. This capability serves as an inspiration for the realization that similar traffic control principles can be harnessed within the realm of LibOS to achieve intelligent resource management. In the context of demanding tasks such as machine learning training that consume vast amounts of resources, as well as in resource-constrained edge devices, meticulous oversight over resource allocation is paramount. This empowers dynamic, autonomous resource allocation, optimizing performance and efficiency across varied workloads and settings, perfectly suiting complex, evolving AI applications and infrastructures.
@@
Microservices architecture heralds a seismic shift in software development paradigms, dismantling centralized structures in favor of agile, decentralized teams, each specializing in a unique microservice[83]. This transition boosts efficiency and reliability, as focused teams oversee discrete services, revolutionizing software engineering approaches. Contrastingly, monolithic systems’ vast scale and tangled dependencies escalate code generation complexities. Automating code in large, interconnected systems demands meticulous integration within a dense module network[84]. LibOS, with their compact, independent nature, simplify code generation, offering fertile ground for efficient automation and AI-driven development.
9.2.2 Multi-agent LLM for OS Development
Previous research has demonstrated the efficacy of employing AI in optimizing, safeguarding, and evaluating system software including operating systems[10][22][28]. Our study reveals a notable dearth of efforts dedicated to leveraging LLMs in the context of operating systems. Nonetheless, we identify at least one distinct domains where LLMs can potentially revolutionize the functioning and capabilities of OSes as Multi-agent LLM for OS Development.
Figure 5: One example for Multi-agent LLM for OS Development
@@ 9.2.2 Multi-agent LLM for OS Development
Recent breakthroughs in LLM-based multi-agent systems, building on the success of single LLMs as unified decision-makers, have unlocked new horizons in complex problem-solving and environment simulation. This progress signals the potential for multi-agent LLMs to revolutionize OS development. Studies like [79] and [80] showcase multi-agent LLM models’ efficacy in software development, from analysis to debugging.
@@
In the realm of OS development, the integration of multi-agent systems fosters a specialized, collaborative, and flexible environment. A constellation of agents, each dedicated to niches such as kernel architecture, driver engineering, or cybersecurity, collectively enhances project productivity[72][73][74][75]. these agents facilitate knowledge exchange and consensus-building akin to human collaboration. They concurrently tackle OS components, recalibrate in response to evolving project dynamics, and oversee integration, testing, and debugging phases, proactively addressing potential issues.Inherently adaptable, the system reassigns agent responsibilities or integrates fresh entities to align with technological advancements, market dynamics, or emerging security vulnerabilities, ensuring continuous OS refinement. Certain agents conduct simulations to assess OS performance under varied scenarios, supplying empirical insights to inform judicious decision-making, preserving the OS’s technological relevance and competitiveness.
9.2.3 LLM AS OS
Future research in the domain of LLMs integrated as OS will focus on following areas.
@@ 9.2.3 LLM AS OS
Enhanced Security Measures
LLMs contribute to OS security by analyzing data for threats, enforcing policies, and managing access. They can predict security incidents, provide real-time code analysis, and offer personalized security guidance, fortifying the system against vulnerabilities.

### v2

@@ 6.4 Analysis
In summary, these advanced OS architectures play a transformative role in advancing AI-OS integration. They cater to the requirements of AI applications, promising to rewrite the rules of computing by blending AI and OS synergies for smarter, more efficient, and secure computing futures.
7 How can we define the different stages of AI-assisted OSs?
Figure 7: Three stages of AI-OS integration
The evolution of AI-OS integration can be delineated into three progressive stages as shown in Figure 7, each marking significant advancements in the integration and utilization of AI within the foundational software that manages computer hardware resources and provides common services for computer programs.
7.1 AI-Powered OS
AI-powered Operating Systems build upon traditional architectures by integrating AI tools to enhance user experiences and optimize resource management efficiency. These enhancements include providing applications with common, system-level intelligent services such as understanding user needs (?), learning habits and preferences, and recommending applications. Additionally, these OSs more efficient and intelligent resource management services that are independent of hardware type and distribution, facilitating cross-device application execution, cross-architecture application migration, and dynamic hardware allocation based on operational data analysis. Furthermore, they support real-time fault detection and repair, dynamic monitoring of system operational states, and real-time adjustment of internal policies to ensure optimal operational performance. Lastly, they leverage AI in software engineering to automate or correct operating system code, thereby improving the efficiency and quality of OS development via automation and intelligence. Currently, most previous jobs focus on this stages.
7.2 AI-Refactored OS
The transition to an AI-Refactored Operating System (OS) involves a profound architectural transformation where intelligent mechanisms are deeply integrated into the OS framework (?, ?, ?). This advancement leads to a sophisticated level of software intelligence, fostering the development of intelligent, modularized OS structures.
This stage paves the way for leveraging AI to conduct a sophisticated analysis of assets derived from open-source operating systems. AI algorithms formulate methodologies for constructing micro-libraries and facilitate the decoupling of existing systems, thereby maximizing the utilization of the current operating system ecosystem. In alignment with the generation requirements of the OS, AI identifies optimal combinations or configurations of these micro-libraries. Additionally, AI enables precise sensing of the operational state of the OS, which supports self-adaptive optimization during runtime, fostering the autonomous growth of the OS.
Moreover, AI-Refactored OS facilitates the strategic planning of coordinated evolutionary pathways among multiple operating systems, promoting a collective process of self-evolution. Ultimately, this approach establishes an ecological paradigm grounded in shared capabilities, positioning itself competitively in the ongoing technological race.
7.3 AI-Driven OS
The AI-Driven OS signifies a radical departure from conventional frameworks by placing AI at the core of OS. It replaces the traditional operational logic with AI-driven logic and employs AI to manage and schedule various computing resources uniformly (?, ?, ?). This phase may see AI wholly or partially supplanting the traditional kernel, replacing heuristic-based decision-making systems with AI-driven operations, and supporting auto-generation and evolution of kernel modules under AI guidance. Additionally, it utilizes AI for comprehensive scheduling of diverse computing resources, achieving unified management and orchestration.
@@ 8 Opportunities for Future Work
During our research,we find that the escalating sophistication of AI workloads, characterized by high-velocity I/O, minimal latency, and bandwidth-intensive requirements, is catalyzing an urgent need for AI-specialized OS optimization (?). As AI permeates diverse sectors, the imperative for a finely calibrated infrastructure escalates, poised to support the burgeoning demands of next-generation applications. The emergence of AI-dedicated hardware has dramatically reshaped AI computation landscapes. These technological marvels, while delivering unrivaled performance, necessitate software ecosystems meticulously attuned to their singular attributes.
In this multifaceted hardware ecosystem (?, ?), the need for an OS that can seamlessly manage and orchestrate AI tasks across a spectrum of devices is more pressing than ever. AI has transcended beyond the realm of traditional neural networks, with innovative models leading the change in natural language processing and graph neural networks adeptly handling intricate relational data. These advanced applications are calling for specialized optimizations that a one-size-fits-all approach cannot provide. Juggling performance, energy conservation, and scalability constitutes a multifaceted challenge, demanding an OS equipped for dynamic adjustment to cater to the singular requisites of every workload.
Heterogeneous adaption for AI
In industrial settings, IoT devices generate massive amounts of data. Deploying deep learning models directly on edge devices (such as sensors, gateways, or edge servers) in equally important as relying solely on cloud-based processing. In a certain sense, deploying on edge devices can be a more challenging scenario because the computational power and bandwidth, of edge devices are more constrained (?, ?, ?). Due to the diversity of edge devices,current AI platforms’ lack of portability across different edge platforms hinders widespread adoption ,and bypassing the kernel and directly accessing hardware while adapting to various hardware configurations can be extremely challenging. Therefore, an operating system that can accommodate multiple hardware types and is specifically tailored for AI applications is a potential solution (?, ?).
@@
Resource-constrained environments, such as edge devices or embedded systems, often operate with tight constraints on memory, processing power, and energy while in scenarios like cloud computing or data centers, minimizing resource consumption directly impacts operational expenses.(?) As OS and applications share the resources, the OS consumption determines the actual operational performance of the algorithmic model. So there is a need to design an efficient OS that can achieve small memory and fast training and inference. There are several promising approaches to achieve this. Apart from bypassing the kernel, unikernel OSs are designed to be fast and lightweight (?, ?, ?).
@@
The problem of privacy preservation in the context of AI is quite different from that in traditional data privacy protection. Increasingly, edge devices are equipped with AI applications. This trend comes with privacy risks as models can leak information about their training data (?). OS need to find ways to provide safeguards for ensuring the confidentiality and integrity of its data and programs.
8.1 Library OS based Intelligent OS
Inspired by previous requirement for OS enhancement for a intelligent OS, we believe Library os is a potential solution for it.
@@ 8.1 Library OS based Intelligent OS
A recently emerging trend of Internet-based software systems is ‘resource adaptive’, thus software systems should be robust and intelligent enough to the changes of heterogeneous resources (?), both physical and logical, provided by their running environment (?). The key principle of such an OS is based on resource disaggregation, resource provisioning as a service, and learning-based resource scheduling and allocation. Meanwhile, the OS should leverage advanced machine/deep learning techniques to derive configurations and policies and automatically learn to tune itself and schedule resources. One general and efficient Spark tuning framework is proposed and applied it to Tencent’s data platform (?).
@@
Microservices excels notably in granular security, isolated updates, and performance optimization through dynamic service management. AI technologies have proven adept at governing microservices, enabling intelligent service orchestration and customization, as highlighted in research (?) and (?), setting the stage for advanced system control and optimization (?). LibOS similarly exhibit gains in agility, composability, enhanced isolation, and facilitated deployment across heterogeneous hardware.
Heterogeneous deployment and minimized OS consumption
In the IoT epoch, a surge in varied endpoint devices underscores the imperative for AI integration. Yet, conventional monolithic OS kernels struggle, weighed down by excessive resource demands and challenges in hardware acclimatization, hindering their agility in today’s rapidly transforming milieu (?). In alignment, LibOS surface as a beacon of hope for versatile, cross-platform deployment amid the AI revolution. The Demikernel project (?) substantiates LibOS’s flexibility, enabling applications to operate fluidly across disparate devices sans customization. In LibOS’s schema, hardware interface components are architected as pluggable modules, amplifying interoperability and malleability.The intrinsic modularity and adaptability simplify AI assimilation into a spectrum of devices, nurturing a more dynamic and reactive IoT framework.
@@
FlexOS (?) presents a modular LibOS framework, comprised of discrete elements that can be segregated using a spectrum of hardware safeguards, complemented by adaptable data sharing protocols and software fortification methods. FlexOS adopts micro-libraries as fundamental units for segmentation. Integration or isolation of these components is dynamically managed, guided by safety and performance requirements.
Drawing from the microservices paradigm’s strengths in isolation and precise access governance, we posit that LibOS offers a strategic avenue to uphold data confidentiality during processing and safeguard overall system integrity. In the AI domain, where security and privacy are paramount(?). LibOS’s robust isolation features make them ideally suited for secure operation in environments with concurrent tasks of varying privacy sensitivity. The innate simplicity and modularity of LibOS create an ecosystem that effectively quarantines critical workloads, bolstering confidence and aligning with rigorous privacy statutes.
@@
A significant advancement brought about by the microservices paradigm is its capacity for orchestration for components, exemplified by tools like ISTIO. This capability serves as an inspiration for the realization that similar traffic control principles can be harnessed within the realm of LibOS to achieve intelligent resource management. In the context of demanding tasks such as ML training that consume vast amounts of resources, as well as in resource-constrained edge devices, meticulous oversight over resource allocation is paramount. This empowers dynamic, autonomous resource allocation, optimizing performance and efficiency across varied workloads and settings, perfectly suiting complex, evolving AI applications and infrastructures.
@@
Microservices architecture heralds a seismic shift in software development paradigms, dismantling centralized structures in favor of agile, decentralized teams, each specializing in a unique microservice (?). This transition boosts efficiency and reliability, as focused teams oversee discrete services, revolutionizing software engineering approaches. Contrastingly, monolithic systems’ vast scale and tangled dependencies escalate code generation complexities. Automating code in large, interconnected systems demands meticulous integration within a dense module network (?). LibOS, with their compact, independent nature, simplify code generation, offering fertile ground for efficient automation and AI-driven development.
Figure 10: One example for Applying Multi-agent LLM to Automate OS Development
@@
Recent breakthroughs in LLM-based multi-agent systems, building on the success of single LLMs as unified decision-makers, have unlocked new horizons in complex problem-solving and environment simulation. This progress signals the potential for multi-agent LLMs to revolutionize OS development. Studies like (?) and (?) showcase multi-agent LLM models’ efficacy in software development, from analysis to debugging.
@@
In the domain of OS development, the integration of multi-agent systems as shown in Figure 10, facilitates a specialized, collaborative, and adaptable environment. A constellation of agents, each dedicated to specific niches such as kernel architecture, driver engineering, or cybersecurity, collectively contributes to enhanced project productivity(?, ?, ?, ?). These agents facilitate knowledge exchange and consensus-building akin to human collaboration. They concurrently tackle OS components, recalibrate in response to evolving project dynamics, and oversee integration, testing, and debugging phases, proactively addressing potential issues.Inherently adaptable, the system reassigns agent responsibilities or integrates fresh entities to align with technological advancements, market dynamics, or emerging security vulnerabilities, ensuring continuous OS refinement. Certain agents conduct simulations to assess OS performance under varied scenarios, supplying empirical insights to inform judicious decision-making, preserving the OS’s technological relevance and competitiveness.
8.2 LLM Based Intelligent OS
Ideas such as those presented in (?) and (?) propose innovative frameworks for current LLM-based operating systems or runtimes. However, in line with our definition of an AI-driven OS, we believe that the generalization capabilities of LLMs position them to function as various sub-modules of operating systems—such as schedulers or filesystems—effectively replacing traditional OS components. This approach can be realized without necessitating changes to existing applications or hardware adaptations, allowing for a smoother transition and integration. By leveraging the adaptive and conversational nature of LLMs, these systems could enhance resource management, task scheduling, and data handling in a more intuitive and efficient manner, ultimately transforming the landscape of operating systems while maintaining compatibility with current software and hardware infrastructures. Future research in the domain of LLMs integrated as OS will focus on following areas.
@@ 8.2 LLM Based Intelligent OS
Instructions Interporation
Traditional operating systems link to applications through system calls, and if large models can accurately translate and understand these instructions, they indeed have the potential to replace the system call layer or even the entire operating system. By leveraging natural language processing and contextual understanding, large models could directly interpret application demands and generate corresponding system-level operations, thereby simplifying the interaction between users and the system. Such a replacement could enhance the system’s flexibility and usability while lowering the technical requirements for developers and users.